OpenStudy (anonymous):

Solve the system using elimination. 5x+4y=12 3x-3y=18

OpenStudy (***[isuru]***):

hi, take 5x + 4y = 12 ----(1) 3x - 3y = 18 ---(2) now multiply (2) by 4 and u will get 12x -12y = 72 ----(3) also... multiply (1) by 3 and u will get.. 15x +12y = 36 ----(4) now add (3) and (4)... ya will find out that "y" will be eliminated ... now solve it for "x".. Did ya get it???

OpenStudy (anonymous):

So would x=-36?

OpenStudy (anonymous):

well i was still doing the equation the other guy showed. sorry. then x=4

OpenStudy (anonymous):

i mean 3

OpenStudy (***[isuru]***):

@radar mmmm.... there is smthing wrong....

OpenStudy (***[isuru]***):

how can u get x - y = 3 by dividing 3x-3y=18 by 3?????

OpenStudy (radar):

Did you try 3 and solve for y, lets see 5(3) +4y =12 15 + 4y = 12 4y = -3 y = - OOOOOOh NOOOOO made error that should be 6 Sorry about that

OpenStudy (anonymous):

im getting so confused

OpenStudy (radar):

@JjJustin the division the 2nd formula, there was an error

OpenStudy (***[isuru]***):

Divide the 2nd equation by 3 getting x - y = 3 is wrong............ u will get x-y = 6

OpenStudy (anonymous):

ok

OpenStudy (radar):

Divide the 2nd equation by 3 getting x - y = 6 Now multiply this by 4 getting 4x -4y = 12 Now add the first equation to eliminate. 4x+4y = 24 ------------ 8x = 36 solve for x and substitute the value of x in one of the original equation and solve for y.

OpenStudy (radar):

Thanks @***[ISURU]*** I made an error which blew the solution.

OpenStudy (anonymous):